Calculate normality and titre of KOH -solution (volume 6.5 ml) if it is 

nessesary for the titration 5.0 ml HCl – solution with CN = 0.1065 mol/l.


1
Expert's answer
2021-02-10T01:44:47-0500

Solution:

The balanced chemical equation:

KOH + HCl → KCl + H2O

According to the equation above:

n(KOH) = n(HCl)

or

CN(KOH) × V(KOH) = CN(HCl) × V(HCl)

Hence,

CN(KOH) = CN(HCl) × V(HCl) / V(KOH)

CN(KOH) = 0.1065 mol L-1 × 5.0 mL / 6.5 mL = 0.08192 mol L-1

CN(KOH) = 0.08192 mol/L


T(KOH) = CN(KOH) × M(KOH) / 1000

The molar mass of KOH is 56.1056 g mol-1.

Hence,

T(KOH) = (0.08192 mol L-1 × 56.1056 g mol-1) / 1000 = 0.004596 g mL-1

T(KOH) = 0.004596 g/mL


Answer:

Normality of KOH solution is 0.08192 mol/L.

Titre of KOH solution is 0.004596 g/mL.